(Related to Checkpoint​ 7.1) ​ (Expected rate of return and risk​)  B. J. Gautney Enterprises is evaluating a security. ​ One-year Treasury bills are currently paying 4.5 percent. Calculate the​ investment's expected return and its standard deviation. Should Gautney invest in this​ security? Probability Return 0.10 negative 4 ​% 0.45 3 ​% 0.35 5 ​% 0.10 10 ​% a.  The​ investment's expected return is nothing​%. ​(Round to two decimal​ places.)
Answer #1

1.
Expected returns=0.1*(-4%)+0.45*3%+0.35*5%+0.10*10%=3.700%

2.
Standard Deviation=sqrt(0.1*(-4%-3.7%)^2+0.45*(3%-3.7%)^2+0.35*(5%-3.7%)^2+0.10*(10%-3.7%)^2)=3.273%

No should not invest..as Treausry bills have zero standard deviation or risk but offer higher return